• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ramesh

    Log entry time 13:06:50 on December17,2007

    Entry number 221650

    keyword=Swing shift (Dec 16) summary

    Shift Summary
    ----------------------

    Left spectrometer:
    Run 22014 ended 16:15 Pb 2.6M 50uA

    Right spectrometer:
    Run 4147 ended 16:15 Pb 2.7M 50uA

    Continuing Part 4 of "Additional Run Plan"
    instructions for the Right Spectrometer.

    Asked MCC for raster size 4x3.5 so that we get 2x2 (as per instructions on whiteboard)

    Run 22015 16:25-16:31 Pb 0.8M 5uA to test raster size (it was 1.5x1.5)
    Asked MCC to set raster size to 4x3.5
    Run 22016 16:35-16:40 Pb 0.6M 5uA raster size is 2.6x2
    Run 22017 16:43-16:48 Pb 0.8M 10uA " " "

    Run 22018 checking raster size
    Raster size still not ok. Asked MCC for 3.2x3.5

    Raster is now 2x2 (spot++)
    ============================

    Run 22019 16:56-17:01 Pb 1M 10uA (raster size ok 2x2)
    Run 22020 17:04-17:11 Pb 1M 5uA
    Run 22021 17:16-17:21 Pb 1.2M 20uA
    Run 22022 17:25-17:30 Pb 1.4M 30uA
    Run 22023 17:33-17:38 Pb 1.4M 40uA
    Run 22024 17:41-17:46 Pb 1.2M 50uA

    17:23 MCC call wants to do spot move (due to hall C). Asked for ~20 min to finish our boiling test.
    17:48 MCC takes beam away. Spot move.

    18:14 beam back, asked for 5uA

    Run 22025 test to set prescales
    Run 22026 18:24-18:29 LH2 .8M 5uA
    Run 22027 18:33-18:38 LH2 .9M 10uA
    Run 22028 18:41-18.43 LH2 1.1M 20uA
    Run 22029 18:50-18:55 LH2 1.3M 30uA
    Run 22030 18:58-19:04 LH2 1.4M 40uA
    Run 22031 19:07-19:12 LH2 1.2M 50uA
    Run 22032 19:15-19:21 LH2 1.2M 60uA
    Run 22033 stopped run beam tripped
    Run 22034 19:33-19:38 LH2 1.3M 70uA

    19:20 Left spectrometer at 23 deg and 735.14 MeV/c (cycled quads)

    Boiling test C12 reference scan:
    ==============================================
    raster 2x2
    ===========

    Left spectrometer:
    --------------------

    Run 4148 20:03-20:09 C12 0.8M 5uA run to test prescales (good run)
    Run 4149 20:17-20:25 C12 1.M 5uA
    Run 4150 20:31-20:38 0.4M ended run, beam off
    Run 4151 stopped beam trips frequently
    Run 4152 20:41-20:47 C12 0.5M 10uA
    Run 4153 20:52-20:57 C12 1.1M 20uA
    Run 4154 21:00-21:05 C12 1.2M 30uA
    Run 4155 21:08-21:13 C12 1.3M 40uA
    Run 4156 21:18-21:23 C12 1.4M 50uA end of run not normal
    Run 4159 22:16-22:22 C12 1.5M 60uA
    Run 4160 22:27-22:32 C12 1.3M 50uA
    Run 4161 22:37-22:46 C12 2.6M 70uA


    20:16 Right spectrometer at 23 deg and 735.14 MeV/c (cycled quads)

    Right spectrometer:
    ---------------------

    Run 22035 20:17-20:25 C12 1.1M 5uA
    Run 22036 20:31-20:38 C12 0.4M 10uA ended run, beam off
    Run 22037 stopped beam trips
    Run 22038 20:41-20:47 C12 0.5M 10uA
    Run 22039 20:52-20:57 C12 1.1M 20uA
    Run 22040 21:00-21:05 C12 1.2M 30uA
    Run 22041 21:08-21:13 C12 1.2M 40uA
    Run 22042 21:18-21:23 C12 1.4M 50uA end of run not normal
    Run 22044 22:16-22:22 C12 1.5M 60uA
    Run 22045 22:27-22:32 C12 1.3M 50uA
    Run 22046 22:37-22:46 C12 2.6M 70uA


    -------------------------------------------------------------------------------------------------------

    21:28 Cannot start a run, tried resetting ROC3, ROC1
    cannot kill coda
    Hall C has similar coda problems
    Problem seems to be sitewide, computer center related (network card?)
    Paged Bob Michaels.

    22:09 Asked for beam. Part of the problem seems to be fixed. STILL CANNOT MAKE HALOG ENTRY
    Lost beam time accounting!

    22:47 Asked for Hall access JP nedds to do target loop 2 work

    23:03 MCC asked for Beam time accounting. Cannot give them any numbers. bta screen is frozen.
    They will use what they have.
    23:07 JP Makis back from Hall, asked for peam permit.
    23:08 Moved target to empty, asked for raster off.
    23:10 Setting both spectrometers as per section #5 of the Additional Run Plan.
    cycling Q2 and Q3 of both spectrometers.
    Momentum set is 738.78 MeV/c for both spectrometers. Angle is 23 deg


    23:25 started target position scan on empty target. Beam current 10uA, raster off [table below]

    23:50 Sandy Philpot called from the computer center: File server has crashed and one vendor person
    is here working on problem. Estimate to fix is of the order of hours!

    BPM A x | BPM B x | rad29_p1 | T1 (Hz) | T3 (Hz) | I_BCM
    --------------------------------------------------------------
    -1.14 | -2.00 | 0.035 | 10 | 19
    -1.03 | -1.50 | 0.030 | 11 | 20
    -0.93 | -1.00 | 0.030 | 11 | 21
    -0.61 | -0.50 | 0.033 | 14 | 23
    -0.054 | 0.004 | 0.036 | 20 | 25
    0.973 | 0.994 | 0.028 | 270 | 350 | 10.26
    0.484 | 0.530 | 0.025 | 25 | 32 | 10.25
    1.369 | 1.600 | 0.037 | 1.5e4 | 2.0e4 | 10.15
    1.377 | 1.516 | 0.037 | 1.5e4 | 2.0e4 ! 10.08
    2.017 | 2.020 | 0.041 | 1.5e4 | 2.0e4 | 10.05
    -3.027 | -3.174 | 0.026 | 13 | 20 | 10.04
    -3.237 | -3.466 | 0.019 | 15 | 25 | 10.01